OpenStudy (anonymous):

find the percent error or the percent of decrease: The circulation of a newsletter decreased from 5200 to 3140. Find the percent of decrease in the circulation to the nearest percent.

OpenStudy (anonymous):

Decrease of value = 5200 - 3140 = 2060 Percentage of decrease = (Decrease of value/ Original Value) x 100% = 2060 / 5200 x 100% = 39.6%